OpenStudy (anonymous):

whats the domain of the function f(x)= log base 2 square root x^2-3

OpenStudy (anonymous):

yes its inside the square root

OpenStudy (anonymous):

\[x^2-3>0\] \[(x-\sqrt{3})(x+\sqrt{3})>0\] \[x<-\sqrt{3}\text { or } x>\sqrt{3}\]

OpenStudy (anonymous):

thank you very much for your help!